# Bound-state confinement after trap-expansion dynamics in integrable systems

### v1

@@ 3 Thermodynamic Bethe Ansatz ($TBA$) for quenches from homogeneous states in the $XXZ$ chain
To derive bound-state confinement after the trap expansion of the $p$-Néel states (cf. (2)), it is necessary to construct the $GGE$ describing the steady state after the quench from the homogeneous $p$-Néel states. Here we provide a self-contained discussion of the derivation of the $GGE$ for a generic pre-quench initial state. We follow Ref. [20, 21].
We employ the framework of the Thermodynamic Bethe Ansatz [2] ($TBA$). The eigenstates of the $XXZ$ chain are identified by a set of complex numbers $\lambda_{j}$, which are called rapidities. Each eigenstate is written as |$|\{\lambda_{j}\}\rangle$, and the rapidities $\lambda_{j}$ label the different quasiparticles in the system. In the thermodynamic limit $L\to\infty$ at fixed density of quasiparticles, it is convenient to work with the rapidity density. Moreover, in the thermodynamic limit generic integrable systems exhibit different families of quasiparticles. They correspond to bound states of elementary, magnon-like, excitations. The corresponding rapidities form string patterns in the complex plane [1]. Then, any thermodynamic macrostate of the system is characterized by a set of rapidity densities $\{\rho_{j}(\lambda)\}_{j=1}^{\mathcal{N}}$, where the number of families ${\mathcal{N}}$ depends on the model under consideration. For the $XXZ$ chain at $\Delta>1$ one has ${\mathcal{N}}\to\infty$, meaning that bound states of arbitrary sized are present. Similar to the densities $\rho_{j}(\lambda)$, one defines the hole densities $\rho_{h,j}(\lambda)$. The particle and hole densities for any legitimate $TBA$ macrostate satisfy the $TBA$ equation [2] as
@@
where the energy $\varepsilon_{n}$ and the quasimomentum $p_{n}$ are defined as
$\varepsilon_{n}(\mu)=\frac{\sinh(\eta)\sinh(n\eta)}{\cosh(n\eta)-\cos(2\mu)},\quad p_{n}(\mu)=-i\sum_{j=1}^{n}\ln\left[\frac{\sin\left(\mu+i\frac{\eta}{2}(n+1-2j)+i\frac{\eta}{2}\right)}{\sin\left(\mu+i\frac{\eta}{2}(n+1-2j)-i\frac{\eta}{2}\right)}\right].$ (9)
Notice that Eq. (3) allows to obtain only one of the $TBA$ densities, either $\rho_{n}(\lambda)$ or $\rho_{h,n}(\lambda)$. Crucially, to uniquely identify a thermodynamic macrostate, one has to determine both the particle and the hole densities. In the following, instead of $\rho_{n}(\lambda)$ and $\rho_{h,n}(\lambda)$ we will often consider the ratios $\eta_{n}(\lambda)$ and $\vartheta_{n}(\lambda)$ defined as

### v2

@@ 3 Thermodynamic Bethe Ansatz ($TBA$) for quenches from homogeneous states in the $XXZ$ chain
To derive bound-state confinement after the trap expansion of the $p$-Néel states (cf. (2)), it is necessary to construct the $GGE$ describing the steady state after the quench from the homogeneous $p$-Néel states. A self-contained discussion of the derivation of the $GGE$ for generic prequench initial states is reported in A.
Here, we employ the framework of the Thermodynamic Bethe Ansatz [2] ($TBA$). The eigenstates of the $XXZ$ chain are identified by a set of complex numbers $\lambda_{j}$, which are called rapidities. Each eigenstate is written as |$|\{\lambda_{j}\}\rangle$, and the rapidities $\lambda_{j}$ label the different quasiparticles in the system. In the thermodynamic limit $L\to\infty$ at fixed density of quasiparticles, it is convenient to work with the rapidity density. Moreover, in the thermodynamic limit generic integrable systems exhibit different families of quasiparticles. They correspond to bound states of elementary, magnon-like, excitations. The corresponding rapidities form string patterns in the complex plane [1]. Then, any thermodynamic macrostate of the system is characterized by a set of rapidity density $\{\rho_{j}(\lambda)\}_{j=1}^{\mathcal{N}}$, where ${\mathcal{N}}$ depends on the model under consideration and where, for $j>1$, $\rho_{j}=\rho_{j}(\mathrm{Re}(\lambda))$ with $\mathrm{Re}(\lambda)$ the center of the string. From now on, $\lambda$ will denote the string center, and we will drop the real part in $\mathrm{Re}(\lambda)$. For the $XXZ$ chain at $\Delta>1$ one has ${\mathcal{N}}\to\infty$, meaning that bound states of arbitrary size are present. Similar to the densities $\rho_{j}(\lambda)$, one defines the hole densities $\rho_{h,j}(\lambda)$. The particle and hole densities for any legitimate $TBA$ macrostate satisfy the $TBA$ equation [2] as
@@
where the bare energy $\varepsilon^{b}_{n}$ and the quasimomentum $p^{b}_{n}$ are defined as
$\varepsilon^{b}_{n}(\mu)=\frac{\sinh(\eta)\sinh(n\eta)}{\cosh(n\eta)-\cos(2\mu)},\quad p^{b}_{n}(\mu)=-i\sum_{j=1}^{n}\ln\left[\frac{\sin\left(\mu+i\frac{\eta}{2}(n+1-2j)+i\frac{\eta}{2}\right)}{\sin\left(\mu+i\frac{\eta}{2}(n+1-2j)-i\frac{\eta}{2}\right)}\right].$ (9)
Notice that Eq. (3) allows to obtain only one of the $TBA$ densities, either $\rho_{n}(\lambda)$ or $\rho_{h,n}(\lambda)$. Crucially, to uniquely identify a thermodynamic macrostate, one has to determine both the particle and the hole densities. The information about the expectation values of local and quasilocal conserved quantities over the initial state allows to fix the $TBA$ densities identifying the $GGE$ that describes the steady state after the quench [25, 26]. In particular, the hole density is fixed by the initial-state expectation values of quasilocal charges $X_{n}(\lambda)$ by the relationship
$\rho_{h,n}(\lambda)=a_{n}-X_{n}\left(\lambda+i\frac{\eta}{2}\right)-X_{n}\left(\lambda-i\frac{\eta}{2}\right),$ (10)
where $a_{n}$ is the same as in  (5) and $\eta$ is defined in  (6). An explicit expression for $X_{n}$ (49) is given in A. Hence, Eq. (3) and Eq. (10) form a system of equations that can be solved to univocally determine $\rho_{n}(\lambda)$ and $\rho_{h,n}(\lambda)$.
In the following, instead of $\rho_{n}(\lambda)$ and $\rho_{h,n}(\lambda)$, we will often consider the ratios $\eta_{n}(\lambda)$ and $\vartheta_{n}(\lambda)$ defined as